### v1

# A Survey of Large Language Models in Psychotherapy: Current Landscape and Future Directions

Abstract
Mental health remains a critical global challenge, with increasing demand for accessible, effective interventions. Large language models (LLMs) offer promising solutions in psychotherapy by enhancing the assessment, diagnosis, and treatment of mental health conditions through dynamic, context-aware interactions. This survey provides a comprehensive overview of the current landscape of LLM applications in psychotherapy, highlighting the roles of LLMs in symptom detection, severity estimation, cognitive assessment, and therapeutic interventions. We present a novel conceptual taxonomy to organize the psychotherapy process into three core components: assessment, diagnosis, and treatment, and examine the challenges and advancements in each area. The survey also addresses key research gaps, including linguistic biases, limited disorder coverage, and underrepresented therapeutic models. Finally, we discuss future directions to integrate LLMs into a holistic, end-to-end psychotherapy framework, addressing the evolving nature of mental health conditions and fostering more inclusive, personalized care.
A Survey of Large Language Models in Psychotherapy:
Current Landscape and Future Directions
Hongbin Nathanks: Equal contribution., Yining Hua11footnotemark: 1, Zimu Wang11footnotemark: 1, Tao Shen, Beibei Yu, Lilin Wang, Wei Wang, John Torous, Ling Chen Australian Artificial Intelligence Institute, University of Technology Sydney Harvard University   Xi’an Jiaotong-Liverpool University   University of Pennsylvania hongbin.na@student.uts.edu.au, yininghua@g.harvard.edu zimu.wang19@student.xjtlu.edu.cn
1 Introduction
Mental health plays an increasingly critical role in current healthcare and social well-being. The high prevalence of common psychological disorders, such as depression and anxiety, has led to a growing demand for accessible and effective psychological interventions. However, the core of psychotherapy resides in dynamic, contextual interpersonal interactions—therapists should continuously assess and adjust their intervention strategies Wampold and Imel (2015) based on the patient’s emotional fluctuations, verbal expressions, and social background, fostering a strong therapeutic alliance Stubbe (2018) to achieve symptom resilience. This deep and flexible process contrasts sharply with traditional NLP, which is typically limited to static or single-task settings.
Figure 1: The dynamic and interrelated network among assessment, diagnosis, and treatment in psychotherapy.
Large language models (LLMs) offer a new perspective to addressing this challenge. By leveraging their capability to model extensive context and perform multi-turn reasoning Wang et al. (2024e); Li et al. (2024b), LLMs can capture rich semantics and emotional signals in dialogues Ma et al. (2025), enabling end-to-end language understanding and generation. In assessment, LLMs can extract potential symptom cues from vague and fragmented expressions Tu et al. (2024); Qiu et al. (2024). During diagnosis, they integrate subjective and objective patient information across multiple utterances Chen et al. (2023a); Ren et al. (2024). In therapeutic interventions, they adapt conversational strategies based on patients’ real-time feedback, enabling more flexible and human-like interactions compared to traditional scripted systems Lee et al. (2024b, d). As a result, LLMs have the potential to surpass the conventional “discrete label recognition” paradigm, evolving toward a model of continuous, progressive clinical reasoning, enabling seamless connections across assessment, diagnosis, and treatment, aligning more closely with therapists’ cognitive process and interaction flow.
[…]
color subtree/.style= fill=#1!10, draw=#1, edge+=draw=darkgray, line width=1pt, for children= color subtree=#1
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=hidden-draw, rounded corners, align=left, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=4.0em,font=,, where level=2text width=5.6em,font=,, where level=3text width=6.2em,font=,, where level=4text width=6.0 em,font=,, [ LLMs in
Psychotherapy [ Assessment, color subtree=AssessmentColor, [ Symptoms [ Detection [ hee So et al. (2024); Tu et al. (2024); Souto et al. (2023); Raihan et al. (2024); Xu et al. (2024);
Gyanendro Singh et al. (2024); Uluslu et al. (2024); Yang et al. (2024a); Mohammadi et al. (2024);
Qiu et al. (2024); Schirmer et al. (2024); Yang et al. (2023) , leaf, text width=26.6em ] ] [ Severity [ Galatzer-Levy et al. (2023); Arcan et al. (2024); Aragon et al. (2024); Skianis et al. (2024);
Wang et al. (2024d) , leaf, text width=26.6em ] ] ] [ Cognition [ Maddela et al. (2023); Qi et al. (2024); Wang et al. (2023); Chen et al. (2023c); Gollapalli et al. (2023); Jiang et al. (2024a)
Lim et al. (2024) , leaf, text width=34.5em ] ] [ Behavior [ Li et al. (2024c); Hoang et al. (2024); Sun et al. (2024); Cohen et al. (2024) , leaf, text width=34.5em ] ] [ Advanced [ Yang et al. (2024b); Wang et al. (2024c); Srivastava et al. (2024) , leaf, text width=34.5em ] ] ] [ Diagnosis, color subtree=DiagnosisColor [ Static Diagnosis [ Galatzer-Levy et al. (2023); Jiang et al. (2024b); Hengle et al. (2024); Lan et al. (2024c); Kuzmin et al. (2024) , leaf, text width=34.5em ] ] [ Dynamic Diagnosis [ Chen et al. (2023a); Lan et al. (2024b); Ren et al. (2024); Lan et al. (2024a) , leaf, text width=34.5em ] ] ] [ Treatment, color subtree=TreatmentColor [ LLM as a Virtual
Therapist [ Xiao et al. (2024); Nie et al. (2024); Lee et al. (2024b); Sharma et al. (2024); Kim et al. (2024);
Lee et al. (2024c, d); Chen et al. (2024) , leaf, text width=34.5em ] ] [ LLM as an Assistive
Tool [ Welivita and Pu (2023); Sharma et al. (2023); Maddela et al. (2023); Moon and Bhattacharyya (2024); Lin et al. (2024);
Na (2024); Brown et al. (2024); Zhan et al. (2024) , leaf, text width=34.5em ] ] [ LLM as Simulated
Patients for Clinician
Education [ Chaszczewicz et al. (2024); Wang et al. (2024b); Yosef et al. (2024); Louie et al. (2024) , leaf, text width=34.5em ] ] [ LLM for Evaluation
and Quality Analysis [ Lee et al. (2024e); Zhang et al. (2024a); Wang et al. (2024a); Chiu et al. (2024); Li et al. (2024a); Shapira and Alfi-Yogev (2024);
Sun et al. (2024); Cohen et al. (2024); Na et al. (2024); Nguyen et al. (2024); Zhang et al. (2024b) , leaf, text width=34.5em ] ] ] ]
Figure 2: Taxonomy of Research on Large Language Models in Psychotherapy.
However, existing research on applying LLMs in this field remains somewhat disjointed. Many studies have utilized LLMs for isolated tasks, such as depression detection Yang et al. (2023); Souto et al. (2023) or diagnosis Jiang et al. (2024b), regarding them as superior feature extractors. Another research line has focused on developing mental health counseling chatbots Chen et al. (2023b); Zhang et al. (2024a); however, these systems remain limited to partial assistance due to insufficient integration with clinical workflows. In other words, although LLMs hold the potential to span the entire continuum from assessment to intervention, they remain limited by the fragmented paradigms of traditional NLP, preventing them from fully leveraging their dynamic, contextual capabilities.
To address these gaps, we introduce the first taxonomy that divides the psychotherapy process into three essential dimensions: Assessment, Diagnosis, and Treatment and provides a systematic review of the recent advancements and challenges of LLMs in each stage. We further examine the current landscape from various perspectives, including the coverage of mental disorders, diversity of linguistic resources, alignment with psychotherapy theories, and the types of techniques employed, thereby sketching the overall distribution and characteristics of existing research. Building on this foundation, we discuss key challenges for the future, including issues of technical coherence, resource and language imbalances, and the disconnect between LLM-based approaches and established psychological practices. Through this comprehensive review and framework, we aim to offer methodological insights to inform future research and facilitate the practical integration of intelligent systems across the entire psychotherapy process.
Organization of This Survey.
We present the first comprehensive survey of recent advancements in applying LLMs to psychotherapy. We introduce a conceptual taxonomy that organizes psychotherapy into three core components—Assessment, Diagnosis, and Treatment—and details their dynamic interrelations (Section §2). We review how LLMs are applied within these components, highlighting their roles in facilitating assessments, refining diagnostic processes, and enhancing treatment strategies (Section §3). We examine current research trends, including symptom and language coverage as well as the distribution of various models and techniques (Section §4). Finally, we discuss open challenges and outline promising directions for future work (Section §5).
2 Conceptual Taxonomy
To establish a standardized framework for understanding psychotherapy, we propose a hierarchical taxonomy aligned with the American Psychological Association’s tripartite model of psychotherapeutic processes [^1] https://www.apa.org/topics/psychotherapy. As illustrated in Figure 1, this taxonomy organizes psychotherapy into three core components: (1) Assessment, (2) Diagnosis, and (3) Treatment, with dynamic interconnections [^2] Throughout this taxonomy, the terms Assessment, Diagnosis, and Treatment specifically refer to the three core components of psychotherapy.. Each component is detailed below.
2.1 Assessment
Definition.
Psychological assessment constitutes the systematic collection and interpretation of data regarding an individual’s cognitive, emotional, and behavioral functioning Cohen et al. (1996); Kaplan and Saccuzzo (2001). This process employs psychometric tests, structured clinical interviews, behavioral observations, and collateral information to establish a multidimensional profile of psychological states Groth-Marnat (2009).
Significance.
As the foundational stage of psychotherapy, assessment provides the empirical basis for understanding a client’s unique psychological landscape. It enables therapists to identify symptom patterns Phillips et al. (2007), track temporal changes Barkham et al. (1993), and contextualize subjective experiences within objective frameworks Groth-Marnat (2009). The continuous nature of psychological assessment allows for real-time adjustments to therapeutic strategies Schiepek et al. (2016), ensuring interventions remain responsive to evolving client needs.
2.2 Diagnosis
Definition.
Diagnosis represents the analytical process of categorizing psychological distress using established nosological systems such as the DSM-5 American Psychiatric Association (2022) and ICD-11 World Health Organization (2019). This involves differentiating normative emotional responses from pathological conditions while considering cultural Teo (2010) and developmental Kawa and Giordano (2012) variables that influence symptom manifestation.
Significance.
Diagnosis serves as the conceptual bridge between assessment and treatment, providing a structured framework for intervention planning Jensen-Doss and Hawley (2011). By aligning clinical observations with standardized criteria, it enhances communication among professionals Craddock and Mynors-Wallis (2014) and facilitates evidence-based decision-making American Psychiatric Association (2006).
2.3 Treatment
Definition.
Treatment includes evidence-based interventions designed to reduce psychological distress and improve functioning American Psychiatric Association (2006). These interventions work by building a therapeutic alliance Elvins and Green (2008), restructuring cognition Ezawa and Hollon (2023), and modifying behavior Martin and Pear (2019), all typically grounded in well-established theoretical orientations.
Significance.
Treatment transforms the theories and information gleaned from assessment and diagnosis into practical interventions Prochaska and Norcross (2018) that directly address the client’s psychological distress Barlow (2021) and foster personal growth Lambert (2013).
2.4 Interrelations
The taxonomy’s components interact through three dynamic processes that define psychotherapy as a complex adaptive system:
Synthesizing (Assessment $\rightarrow$ Diagnosis)
The dialectical integration of observational data with nosological frameworks enables diagnostic classifications to contextualize assessment findings, synthesizing the patient’s various symptoms and behavioral patterns into a diagnostic result Rencic et al. (2016).
Framing (Diagnosis $\rightarrow$ Treatment)
Diagnosis functions as a framing mechanism, integrating complex and diverse symptoms into a coherent classification that establishes a clear blueprint for treatment American Psychiatric Association (2022).
Customization (Assessment $\rightarrow$ Treatment)
A process where treatment plans are continuously refined based on assessment results, considering individual differences without being constrained by diagnostic labels, to enhance therapeutic effectivenesss Waszczuk et al. (2017).
3 LLMs in Psychotherapy
3.1 Assessment
Study Text Granularity Best Technique NLP Task Assessment Focus
Symptom Detection
Yang et al. (2023) Single Post Emotion Prompting BC/MCC/EG Multiple Symptoms
hee So et al. (2024) Multi-turn Dialogue Fine-Tuning MLC/IE/SUM Multiple Symptoms
Tu et al. (2024) Multi-turn Dialogue Few-Shot Prompting MLC/IE/SUM PTSD
Souto et al. (2023) Single Post Fine-Tuning MLC/EG Depression
Raihan et al. (2024) Single Post Few-Shot Prompting MCC Multiple Symptoms
Gyanendro Singh et al. (2024) Posts From One User Chain-of-Thought IE/SUM Suicidal Ideation
Uluslu et al. (2024) Posts From One User Role Prompting IE/SUM Suicidal Ideation
Yang et al. (2024a) Single Post Fine-Tuning BC/MCC/EG Multiple Symptoms
Xu et al. (2024) Single Post Fine-Tuning BC/EG Multiple Symptoms
Mohammadi et al. (2024) Single Post Few-Shot Prompting MLC Multiple Symptoms
Qiu et al. (2024) Single Post Fine-Tuning MLC Suicidal Ideation
Schirmer et al. (2024) Single Post Zero-Shot Prompting BC PTSD
Symptom Severity
Galatzer-Levy et al. (2023) Multi-turn Dialogue Zero-Shot Prompting TR Depression/PTSD
Arcan et al. (2024) Multi-turn Dialogue Zero-Shot Prompting TR Depression/Anxiety
Aragon et al. (2024) Posts From One User Zero-Shot Prompting TR Depression
Wang et al. (2024d) Posts From One User Zero-Shot Prompting TR Depression
Skianis et al. (2024) Single Post Zero-Shot Prompting TR/MCC Depression/Suicide
Cognition
Maddela et al. (2023) Single Sentence Few-Shot Prompting MLC Cognitive Distortions
Qi et al. (2024) Single Post Fine-Tuning MLC Cognitive Distortions
Wang et al. (2023) Single Sentence Few-Shot Prompting MCC Cognitive Distortions
Chen et al. (2023c) Single-turn Dialogue Zero-Shot Prompting BC/MCC/EG Cognitive Distortions
Gollapalli et al. (2023) Single Post Zero-Shot Prompting MLC Maladaptive Schemas
Jiang et al. (2024a) Single Post Zero-Shot Prompting MCC/SUM Cognitive Pathways
Lim et al. (2024) Single-turn Dialogue Multi-Agent Debate MCC Cognitive Distortions
Behavior
Li et al. (2024c) Single Post Zero-Shot Prompting MLC/EG Interpersonal Risk
Hoang et al. (2024) Sentence From Dialogue Few-Shot Prompting MCC MI-Adherent Behaviors
Sun et al. (2024) Sentence From Dialogue Zero-Shot Prompting MCC MI-Adherent Behaviors
Cohen et al. (2024) Sentence From Dialogue Zero-Shot Prompting MCC MI-Adherent Behaviors
Table 1: Comparison of Psychological Assessment Studies by Input Characteristics and Methodology. MLC: Multi-Label Classification, IE: Information Extraction, SUM: Summarization, MCC: Multi-Class Classification, BC: Binary Classification, TR: Text Regression, EG: Explanation Generation. Studies are categorized through text granularity, optimal technical approach (Best Technique), NLP task formulation, and specific assessment focus.
Symptom Detection
leverages LLMs to identify mental health conditions including depression, anxiety, PTSD, and suicidal ideation, demonstrating robust performance and multidimensional applicability across diverse scenarios. Yang et al. (2023) systematically evaluated GPT-3.5, InstructGPT3, and LLaMA models across 11 datasets, revealing that emotion-enhanced chain-of-thought prompting improves interpretability yet remains inferior to specialized supervised methods. hee So et al. (2024) achieved 70.8% zero-shot symptom retrieval accuracy in Korean psychiatric interviews using GPT-4 Turbo, while their fine-tuned GPT-3.5 attained 0.817 multi-label classification accuracy. Clinical applications show particular promise, as Tu et al. (2024) leveraged GPT-4 and Llama-2 to automate PTSD assessments through information extraction from 411 interviews, significantly enhancing diagnostic practicality.
Social media analysis benefits from approaches like Souto et al. (2023)’s interpretable depression detection framework, which demonstrated strong performance across Vicuna-13B and GPT-3.5 environments. Resource development advances include Raihan et al. (2024)’s MentalHelp dataset with 14 million instances, validated through GPT-3.5 zero-shot evaluations. For suicidal ideation monitoring, Gyanendro Singh et al. (2024) and Uluslu et al. (2024) achieved state-of-the-art evidence extraction in the CLPsych 2024 shared task through innovative prompting strategies. Open-source initiatives like MentaLLaMA by Yang et al. (2024a) and Mental-LLM by Xu et al. (2024) enable multi-symptom detection via instruction-tuned LLaMA variants, though Mohammadi et al. (2024)’s WellDunn framework reveals persistent gaps in GPT-family models’ explanation consistency.
Cross-lingual adaptations include Qiu et al. (2024)’s PsyGUARD system based on fine-tuned CHATGLM2-6B for Chinese suicide risk assessment, while Schirmer et al. (2024) demonstrated domain-specific RoBERTa models outperforming GPT-4 in cross-domain PTSD pattern analysis, highlighting the critical balance between model specialization and interpretability.
Symptom Severity
focuses on estimating the level of mental health condition intensity, particularly for depression, anxiety, and PTSD. Clinical evaluations reveal Med-PaLM 2’s zero-shot depression scoring attains clinician-level alignment on interview data (Galatzer-Levy et al., 2023), though with limited PTSD generalizability. When benchmarked against specialized Transformers on DAIC-WOZ dataset Gratch et al. (2014), ChatGPT and Llama-2 exhibit moderate efficacy (Arcan et al., 2024), suggesting domain-specific architectures retain advantages in structured assessments. Shifting attention to social media data, Aragon et al. (2024) proposed a pipeline that retrieves depression-relevant text, summarizes it according to the Beck Depression Inventory (BDI) Jackson-Koku (2016), and then utilizes LLMs to predict symptom severity, achieving performance similar to expert evaluations on certain measures. In a similar vein, Wang et al. (2024d) introduced an explainable depression detection system that leverages multiple open-source LLMs to generate BDI-based answers, reporting near state-of-the-art performance without additional training data. Cross-lingual extensions emerge through Skianis et al. (2024)’s framework enabling severity prediction across 6 languages and 2 mental conditions.
Cognition
centers on identifying and understanding maladaptive thinking patterns, such as cognitive distortions and early maladaptive schemas, using LLMs. Maddela et al. (2023) introduced a cognitive distortion dataset and employed a few-shot strategy with GPT-3.5 to generate, classify, and reframe them, while Qi et al. (2024) constructed two Chinese social media benchmarks for cognitive distortion detection and suicidal risk assessment, demonstrating that fine-tuned LLMs are more closely than zero-/few-shot methods to supervised baselines. In a related effort, Wang et al. (2023) released the C2D2 dataset containing 7,500 Chinese sentences with distorted thinking patterns. Expanding on detection methods, Chen et al. (2023c) proposed a Diagnosis of Thought prompting approach for GPT-4 and ChatGPT, which breaks down patient utterances into factual versus subjective content and supports the generation of interpretable diagnostic reasoning. Beyond cognitive distortions, Gollapalli et al. (2023) investigated zero-shot approaches with GPT-3.5 to identify early maladaptive schemas in mental health forums, highlighting challenges in label interpretability and prompt sensitivity. Complementarily, Jiang et al. (2024a) presented a hierarchical classification and summarization pipeline to extract cognitive pathways from Chinese social media text, underscoring GPT-4’s strong performance albeit with occasional hallucinations. Finally, Lim et al. (2024) introduced a multi-agent debate framework for cognitive distortion classification, reporting substantial gains in both accuracy and specificity by synthesizing multiple LLM opinions before forming a final verdict.
Behavior
highlights how user actions—or in the case of Motivational Interviewing (MI), language itself—can serve as a measurable indicator of one’s readiness for change. For instance, Li et al. (2024c) introduced the MAIMS framework, employing mental scales in a zero-shot setting to identify interpersonal risk factors on social media, thereby enhancing both interpretability and accuracy. In clinical dialogues, Hoang et al. (2024) demonstrated how LLMs can automatically detect a client’s motivational direction (e.g., change versus sustain talk) and commitment level, offering valuable insights for MI-based interventions. Extending such analyses to bilingual settings, Sun et al. (2024) proposed the BiMISC dataset and prompt strategies that enable LLMs to code MI behaviors across multiple languages with expert-level performance. Lastly, Cohen et al. (2024) presented MI-TAGS for automated annotation of global MI scores, illustrating how context-sensitive modeling can approximate human annotations in psychotherapy transcripts.
Advanced
research has evolved beyond foundational assessment tasks to emphasize novel methodological paradigms, bias mitigation, and domain-specific summarization frameworks. For instance, Yang et al. (2024b) introduced PsychoGAT—an interactive, game-based approach that transforms standardized psychometric instruments into engaging narrative experiences, improving psychometric reliability, construct validity, and user satisfaction when measuring constructs such as depression, cognitive distortions, and personality traits. In parallel, Wang et al. (2024c) systematically investigated potential biases in various LLMs across multiple mental health datasets, revealing that even high-performing models exhibit unfairness related to demographic factors. The authors proposed fairness-aware prompts to substantially reduce such biases without sacrificing predictive accuracy. Furthermore, Srivastava et al. (2024) presented the PIECE framework, which adopts a planning-based approach to domain-aligned counseling summarization, structuring and filtering conversation content before integrating domain knowledge.
3.2 Diagnosis
Static Diagnosis
is based on a fixed set of data, typically derived from complete dialogues or social media posts. Galatzer-Levy et al. (2023) highlighted the effectiveness of Med-PaLM 2 in psychiatric condition assessment from patient interviews and clinical descriptions without specialized training. Similarly, Jiang et al. (2024b) showcased LLMs’ superior performance on depression and anxiety detection on Russian datasets, particularly with noisy or small datasets. Hengle et al. (2024) evaluated PLMs and LLMs on multi-label classification in depression and anxiety, underscoring the ongoing challenges in applying LMs to mental health diagnostics. Besides, Lan et al. (2024c) introduced DORIS, a depression detection system integrating text embeddings with LLMs, utilizing symptom features, post-history, and mood course representations to make diagnostic predictions and generate explanatory outputs. Kuzmin et al. (2024) developed ADOS-Copilot for ASD diagnosis through diagnostic dialogues, employing In-context Enhancement, Interpretability Augmentation, and Adaptive Fusion based on real-world ADOS-2 clinic scenarios.
Dynamic Diagnosis
involves real-time evaluation based on ongoing, interactive conversations between the patient and LLM, enabling more personalized and contextually relevant insights. Chen et al. (2023a) simulated psychiatrist-patient interactions with ChatGPT, in which the doctor chatbot focused on role, tasks, empathy, and questioning strategies, while the patient chatbot emphasized symptoms, language style, emotions, and resistance behaviors. Lan et al. (2024b) introduced the Symptom-related and Empathy-related Ontology (SEO), grounded in DSM-5 and Helping Skills Theory, for depression diagnosis dialogues. Ren et al. (2024) dissected the doctor-patient relationship into psychologist’s empathy and proactive guidance and introduced WundtGPT that integrated these elements. Lan et al. (2024a) further presented the AMC, a self-improving conversational agent system for depression diagnosis through simulated dialogues between patient and psychiatrist agents.
3.3 Treatment
LLM as a Virtual Therapist
centers on leveraging LLMs to directly engage in therapeutic conversations, often adopting multi-turn dialogues that incorporate recognized psychotherapeutic frameworks. For instance, Xiao et al. (2024) proposed HealMe to facilitate cognitive reframing and empathetic support in line with established psychotherapy principles. Likewise, Nie et al. (2024) introduced CaiTI, a system embedded in everyday smart devices that conducts assessments of users’ daily functioning and delivers psychotherapeutic interventions through adaptive dialogue flows. In a similar vein, Lee et al. (2024b) presented CoCoA, specializing in identifying and resolving cognitive distortions via dynamic memory mechanisms and CBT-based strategies, while Sharma et al. (2024) proposed a step-by-step approach guiding users to execute self-guided cognitive restructuring through multiple interactive sessions. Beyond standard CBT protocols, Kim et al. (2024) focused on aiding psychiatric patients in journaling their experiences, thereby offering richer clinical insights, whereas Lee et al. (2024c) developed a multi-round CBT dataset to refine LLMs for direct counseling-like interactions. Additionally, multi-agent frameworks like MentalAgora (Lee et al., 2024d) highlighted personalized mental health support by integrating multiple specialized agents, and Chen et al. (2024) further explored “mixed chain-of-psychotherapies” to combine various therapeutic methods, aiming to enhance the emotional support and customization delivered by chatbot interactions.
LLM as an Assistive Tool
refrains from providing a holistic therapy role but instead offers targeted support such as rewriting suboptimal counselor responses, generating controlled reappraisal prompts, or aiding clinicians in specific tasks. For example, Welivita and Pu (2023) proposed to rewrite responses that violate MI principles into MI-adherent forms, ensuring more consistent therapeutic dialogue. Meanwhile, Sharma et al. (2023) and Maddela et al. (2023) focused on generating single-turn reframes of negative thoughts—often anchored in cognitive distortions—through controlled language attributes. On the detection side, Moon and Bhattacharyya (2024) built a multimodal pipeline to identify depression and provide CBT-style replies, albeit with an emphasis on technological assistance rather than full-fledged therapy. In the Chinese context, Lin et al. (2024) combined cognitive distortion detection with “positive reconstruction,” demonstrating a single-round rewrite approach for negative or distorted statements, while Na (2024) showcased a structured Q&A format that offers professional yet succinct CBT-based responses. From a knowledge-distillation angle, Brown et al. (2024) demonstrated how smaller models could replicate GPT-4’s MI-style reflective statements, and Zhan et al. (2024) introduced a lighter-weight framework RESORT to guide smaller LLMs toward effective cognitive reappraisal prompts, thus enabling broader accessibility of self-help tools.
LLM as Simulated Patients for Clinician Education
pivots toward generating synthetic yet realistic patient behaviors or multi-level feedback to train or support mental health practitioners. For instance, Chaszczewicz et al. (2024) leveraged LLMs to deliver multi-tier feedback on novice peer counselors’ conversational skills, significantly reducing the need for continuous expert oversight. Similarly, Wang et al. (2024b) using LLM-driven patient simulations that help trainees practice CBT core skills in a controlled, repeatable setup. In the realm of assessing therapy quality, Yosef et al. (2024) showcased a digital patient system to evaluate MI sessions, employing AI-generated transcripts to differentiate novice, intermediate, and expert therapeutic skill levels. Complementarily, Louie et al. (2024) offered Roleplay-doh, a pipeline wherein domain experts craft specialized principles that guide LLM-based role-playing agents, thereby providing customizable training for new therapists.
LLM for Evaluation and Quality Analysis
targets the appraisal of therapy dialogue, counselor techniques, and treatment processes, typically without delivering direct interventions to clients. For instance, Lee et al. (2024e) augmented crisis counseling outcome prediction by fusing annotated counseling strategies with LLM-derived features, achieving substantially improved accuracy. In the Chinese context, Zhang et al. (2024a) introduced CPsyCoun, employing reports-based dialogue reconstruction and automated evaluation to verify counseling realism and professionalism. Beyond single-session analyses, Wang et al. (2024a) used simulated clients to assess perceived therapy outcomes, while Chiu et al. (2024) created the BOLT framework for systematically comparing LLM-based therapy behaviors with high- and low-quality human sessions. Further extending to online counseling, Li et al. (2024a) proposed an LLM-based approach to measure therapeutic alliance, whereas Shapira and Alfi-Yogev (2024) delineated therapist self-disclosure classification as a new NLP task. In the MI domain, Sun et al. (2024) and Cohen et al. (2024) collected bilingual transcripts to systematically annotate therapist–client exchanges for behavior coding and global scores, respectively. Additionally, multi-session perspectives emerge in Na et al. (2024), who proposed IPAEval to track long-term progress from the client’s viewpoint, and Nguyen et al. (2024) analyzed conversation redirection and its impact on patient–therapist alliance over multiple sessions. Finally, Iftikhar et al. (2024) and Zhang et al. (2024b) explored the disparities between LLM- and human-led CBT sessions, highlighting gaps such as empathy and cultural nuance while also introducing CBT-Bench to probe LLMs’ deeper psychotherapeutic competencies.
4 Current Landscape
Figure 3: Distribution analysis of the current landscape.
Our survey encompasses a total of 69 studies in the field of LLMs in psychotherapy. Specifically, 33 studies address assessment, 9 focus on diagnosis, and 32 concentrate on treatment, with 5 studies overlapping across these dimensions. Approximately 74% of the studies employed commercial large language models, while about 77% used prompt-based techniques. This distribution highlights an imbalance in research focus across different stages of the psychotherapy process and reflects a heavy reliance on commercial models and prompt technologies.
Figure 3 presents a comprehensive analysis of the current research landscape in this field. Panel (a) reveals a significant linguistic bias in existing studies, with English-language corpora dominates. While there are limited studies involving Korean and Dutch languages, this highlights a substantial gap in multilingual research approaches. Panel (b) quantitatively demonstrates the distribution of mental health research focuses. Mental disorder-related studies constitute 32% of the total research corpus (represented by the orange outer ring). Within this subset, depression-focused research accounts for 50% of mental disorder studies, followed by anxiety-related research. This distribution indicates a concerning imbalance, where common conditions receive disproportionate attention while more complex disorders, such as bipolar disorder, remain understudied. The analysis of psychotherapy theories in panel (c) uncovers another critical gap in the field. Only 32.8% of the studies incorporate psychotherapy theories in their methodological approach. Notably, emerging therapeutic frameworks, such as humanistic therapy, are particularly underrepresented in current research applications.
5 Future Directions
Integrative Psychotherapy Framework.
While many existing studies focus on a single dimension of psychotherapy, real-world practice involves a continuous process that spans assessment, diagnosis, and intervention Waszczuk et al. (2017). Moreover, these stages typically unfold over multiple sessions, necessitating iterative, multi-turn interactions that incorporate the evolving context of each patient. Future work could therefore aim to develop an end-to-end conversational framework that seamlessly spans from initial evaluation to personalized intervention. By maintaining a system grounded on ongoing, context-sensitive engagement, models could dynamically update assessments and diagnoses over time, ultimately providing more responsive and individualized care.
Addressing Evolving and Multifaceted Nature of Psychotherapy.
Psychotherapy commonly involves shifting symptoms, comorbidities, and nuanced patient experiences, making static or single-label predictions insufficient. Models should integrate multi-label and temporal data to capture how symptoms and emotional states evolve, while avoiding the pitfalls of incomplete symptom detection. For instance, focusing solely on the depressive features of a bipolar patient could lead to an inaccurate diagnosis if the manic phase is overlooked Lee et al. (2024a). Furthermore, current research suggests that LLMs often struggle with multi-label tasks hee So et al. (2024); Mohammadi et al. (2024), highlighting the need for improved model architectures and algorithms that can better account for these complexities.
Resource Infrastructure and Open-Source Tools.
Current research heavily relies on commercial closed-source models, lacking reproducible open-source evaluation methods and multilingual data. Notably, developing multilingual datasets should not solely rely on translating English resources, as psychological research indicates that cultural context plays a critical role in mental health. English-based translations cannot fully substitute for culturally specific data from other languages Watters (2010); Abdelkadir et al. (2024).
Broadening Scope of Disorders and Therapeutic Approaches.
Most studies to date have concentrated on prevalent conditions such as depression and anxiety, leaving complex or less common disorders underexplored. Additionally, research tends to focus on a limited range of therapeutic modalities—primarily cognitive behavioral therapy. Future work could broaden both the range of disorders and the variety of therapeutic approaches, such as humanistic Schneider and Krug (2010) and dialectical behavior therapy Lynch et al. (2006), to better reflect clinical realities Norcross et al. (2022). Such an expansion could deepen the theoretical underpinnings of LLM-based psychotherapy tools and enhance the quality and relevance of digital interventions.
6 Conclusion
LLMs hold significant promise for revolutionizing psychotherapy by enhancing assessment, diagnosis, and treatment processes through dynamic, context-sensitive interactions. Despite the progress made, key challenges such as linguistic biases, limited disorder coverage, and underrepresented therapeutic models persist. Future research should focus on creating integrative, multi-turn systems that span the entire psychotherapy process while addressing the evolving nature of mental health conditions. Expanding resources, embracing diverse therapeutic approaches, and improving model architectures will be crucial in making LLM-driven psychotherapy tools more effective, inclusive, and adaptable.
Limitations
This survey paper, while comprehensive for LLMs in psychotherapy, has several limitations: 1) The studies reviewed primarily focus on the application of LLMs in psychotherapy, and there may be relevant research in adjacent fields or interdisciplinary domains that was not included. 2) Due to the rapidly evolving nature of this area, some recent advancements may not be captured. The scope of this survey is limited to the available literature and may overlook emerging trends or unpublished findings. 3) The review primarily examines studies in English, which could introduce a bias towards research from English-speaking countries, potentially overlooking important cultural perspectives. 4) While we provide a taxonomy of LLM applications in psychotherapy, this framework may not fully encompass the complexity of real-world clinical settings or the diverse range of therapeutic approaches currently in practice.